# SDFs from Unoriented Point Clouds using Neural Variational Heat Distances

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: We compute neural SDFs from unoriented point clouds (left) by first computing a small time step of heat flow (middle) and then using its gradient directions to solve for a neural SDF (right).
@@ 1 Introduction
We propose to lift the heat method to the neural setting (see fig. 1). To that end, we define two well-posed variational problems, which on the one hand, can be optimized well with stochastic gradient descent using a standard network architecture, and on the other hand, have existence guarantees. As in the heat method, in the first step we solve for a time step of the heat flow and compute using this the gradient of the unsigned distance function, exploiting the relation between the heat kernel and the distance. Unlike in the heat method, we do not have a discretization of the domain, as we work with neural representations. Thus, to compute the gradients, we use a variational formulation for a backward Euler timestep of the heat flow [31] in the spirit of the general class of minimizing movement schemes[26], see fig. 2 (center).
@@
In the second step, we fit the computed gradients to obtain the signed distance, taking into consideration that the input point cloud is unoriented, leading to accurate level sets in a narrow band near the surface; see fig. 2 (right).
@@ 1.1 Related Work
The seminal paper SIREN [30] uses an eikonal loss, which combined with periodic activation functions, leads to mostly good results. As the results are very initialization dependent, DiGS [2] has proposed a geometric initialization that improves the result for unoriented point clouds, as well as an additional regularization with a divergence-based loss. StEik [35] has discussed the instabilities inherent in the eikonal equation and proposed additional regularizing losses, as well as using a quadratic neural network instead of the usual linear one. Concurrently with our work, HotSpot [33], building on the quadratic network of StEik, has proposed adding a regularizing loss that encourages the exponent of the SDF to fulfill the heat equation. This approach, like ours, incorporates the relation between the distance function and the heat equation. However, their formulation is still based on the eikonal equation, making it non-convex and more difficult to optimize, whereas we propose to solve two convex functionals one after the other. Furthermore, the density of the point cloud is not taken into account in their method, leading to problematic regions for non-uniform point cloud sampling.
@@
Minimizing movements schemes are a robust and effective time discretization for gradient flow type evolution problems in metric spaces [10]. In [26], deep learning-based minimizing movement schemes for a variety of PDEs of gradient flow type are proposed. In [16], a structure-preserving Eulerian neural network discretization for the numerical solving of $L^{2}$-gradient flows is presented. These schemes are based on the variational principle of the energy-dissipation law, which ensures the monotonic decay of the free energy and avoids unphysical states.
We take inspiration from these machine learning approaches to properly capture the gradient flow dynamics for the heat equation in Euclidean and Riemannian spaces.
@@
Using the heat equation for computing approximate geodesic distances has been introduced to geometry processing by Crane et al. [8, 9]. This method is based on two main steps. First, diffusing heat for a short time step starting from heat concentrated at the source points. Then, the normalized gradients of the heat function are used to compute the approximate geodesic distances by solving a Poisson equation. Many generalizations and improvements have been proposed, the latest being generalization to signed distance functions (GSD) [15] by diffusing signed normals. To the best of our knowledge, this concept has not been used in the neural setting so far. As the original heat method, we also diffuse a scalar function from the input sources. However, we incorporate additional constraints in the second step, which allows us to correctly orient the heat gradients. Hence, unlike GSD, we start with unoriented point clouds and compute a signed distance field.
@@
  • compute accurate neural SDFs for spatially varying point cloud densities,
  • demonstrate state-of-the-art surface reconstruction and accurate recovery of gradients of the SDF,
@@ 2 Method
Throughout this paper, we consider a surface $\mathcal{S}$ which is the boundary of a Lipschitz set, contained in $[-1,1]^{3}$ such that the complement of this set is connected. Later, we will assume additional smoothness of $\mathcal{S}$. Following the original heat method [8], our method comprises two steps. First (Sec. 2.1), we compute the solution $u\colon\mathbb{R}^{3}\to\mathbb{R}$ of a small time step of the heat equation with the mean surface measure as initial data. This measure is approximated via a weighted sum of point measures of a not necessarily uniformly distributed input point cloud. In the second step (Sec. 2.2), we compute the SDF $\phi\colon\mathbb{R}^{3}\to\mathbb{R}$, by requiring that (1) $|\phi|$ is small on the set of input points, (2) the direction of $\nabla\phi$ is aligned with the direction of $\nabla u$. To allow unoriented point clouds as input, we automatically find regions $\mathcal{B}^{+}$ and $\mathcal{B}^{-}$ on the two sides of the cloud, and encourage $\phi$ to have the correct sign in these regions. Both $u$ and $\phi$ are parameterized with a neural network (see Sec. 3 for the details). The regions $\mathcal{B}^{\pm}$ are computed using a simple deterministic algorithm.
2.1 (Unoriented) Point Cloud to (unoriented) normals
In the first step, we use that the normalized gradients of a short-time solution of the heat equation on $\Omega$, with the mean surface area density of a smooth surface as initial data, approximate the gradient of the distance function ${\mathrm{dist}\,}(\cdot,\mathcal{S})$ in the vicinity of the surface (cf. Crane et al [8] for the case on manifolds). We restrict here to a computational domain $\Omega\coloneqq(-1.2,1.2)^{3}$ where the surface $\mathcal{S}$ is scaled to fit into $[-1,1]^{3}$.. In the simple case, for initial data $u^{0}\in L^{2}(\Omega)$ a single timestep $\tau$ of the heat equation
@@
with $u^{\tau}\approx u(\tau,\cdot)$. As we want to parametrize the solution $u$ using a neural network, we prefer a variational approach to compute the discrete time step of heat flow instead of a least-squares loss for the residual of 1, as this would increase the condition. Hence, we turn to the simple minimizing movements approach [10] (cf. [26] in the context of neural networks)
$u^{\tau}\coloneqq\operatorname*{arg\,min}_{u}\left({\mathcal{E}_{\mathrm{MM}}}(u)\coloneqq\int_{\Omega}(u-u^{0})^{2}+\tau|\nabla u|^{2}\,\mathrm{d}x\right).$ (3)
The corresponding Euler-Lagrange equation reads as
$0=2\int_{\Omega}(u^{\tau}-u^{0})\vartheta+\tau\nabla u^{\tau}\cdot\nabla\vartheta\,\mathrm{d}x$
for all test functions $\vartheta\in H^{1}(\Omega)$, which is the weak formulation of (2). Here, $H^{1}(\Omega)$ is the space of $L^{2}$ functions on $\Omega$ with weak derivatives also in $L^{2}$.
The case of interest in this paper is the mean two-dimensional surface measure as initial data. The application of this measure to a function $\phi$ is defined as $\fint_{\mathcal{S}}\phi\,\mathrm{d}a\coloneqq\left(\int_{\mathcal{S}}\,\mathrm{d}a\right)^{-1}\int_{\mathcal{S}}\phi\,\mathrm{d}a$.
If $\mathcal{S}$ is the boundary of a Lipschitz set, this measure is well-defined on $L^{2}$ functions on $\mathcal{S}$ and thus by the trace theorem on $H^{1}(\Omega)$. For this initial data, the generalization of the minimizing movement scheme (3) reads as
$u^{\tau}\!\!\coloneqq\!\operatorname*{arg\,min}_{u}\!\left(\!{\mathcal{E}_{\mathrm{MM}}}(u)\!\coloneqq\!\!\int_{\Omega}\!\!u^{2}+\tau|\nabla u|^{2}\,\mathrm{d}x\!-\!2\fint_{\mathcal{S}}\!u\,\mathrm{d}a\!\right)$ (4)
and the associated Euler-Lagrange equation for $u^{\tau}$ is given by
@@
for all $\vartheta\in H^{1}(\Omega)$, which is the adapted weak form of (2). Note that the counterpart of the term $\int_{\Omega}(u^{0})^{2}\,\mathrm{d}x$ in (3) is independent of $u$, thus irrelevant for the optimization. The following proposition provides the existence of the discrete heat solution $u^{\tau}$.
@@
Under the above assumptions there exists a unique minimizer $u^{\tau}$ on $H^{1}(\Omega)$, which solves $0=\partial_{u}{\mathcal{E}_{\mathrm{MM}}}(u)$.
The proof can be found in Appendix A.
@@
which we assume to be well-defined almost everywhere, yields for smooth surfaces $\mathcal{S}$ a good approximation for the gradient of the distance function, i.e. the unoriented normal field of $\mathcal{S}$, which is well-defined up to the medial axis. However, note that this gradient is consistent only in the near field of the surface and corresponds to the unsigned distance and not the signed we actually would like to compute (fig. 3).
@@
2.2 (Unoriented) Normals to SDF
In the second step, we assume that we are given an approximate, unoriented normal field $n^{\tau}$ and aim for the SDF $\phi$ of the surface. To this end, we pick up the ansatz by Crane et al. [8] and consider a modification of the second step of their heat method. We construct an SDF approximation from $n^{\tau}$ with $\phi<0$ inside the surface, where the gradient should point in the opposite direction of $n^{\tau}$, and conversely, $\phi>0$ outside the surface with the gradient pointing in the same direction as $n^{\tau}$ (fig. 4). To this end, we seek a minimizer of the normal fitting objective
@@
where $[\phi<0]=\{x\in\Omega\mid\phi(x)<0\}$ and $[\phi\geq 0]=\{x\in\Omega\mid\phi(x)\geq 0\}$. Here, to have a decomposition of the domain, we arbitrarily included the set $\{\phi=0\}$ in the second term, which does not make a difference since this set is expected to be lower dimensional. Furthermore, we want the zero-level set of $\phi$ to coincide with the given surface, i.e. we take into account the surface fitting objective
@@
The sum ${\mathcal{E}_{n}}+{\mathcal{E}_{\mathrm{fit}}}$ has at least four minima: the signed distance, the unsigned distance, and their respective negatives. Hence, we need an additional term to favor the SDF as the desired solution. For this, we assume to be given sets $\mathcal{B}^{-}$ on the inside of $\mathcal{S}$ and $\mathcal{B}^{+}$ on the outside, cf. section 3.2.1 on how to generate these sets automatically. Now, we define the orientation objective
@@
which promotes that $\phi\leq 0$ in $\mathcal{B}^{-}$ and $\phi\geq 0$ in $\mathcal{B}^{+}$. Finally, our overall objective for computing the SDF becomes
@@
In fact, it will turn out that, in our descent scheme, ${\mathcal{E}_{\mathcal{B}}}$ vanishes after a few descent iterations, which allows us to choose $\lambda_{\mathcal{B}}=1$. We obtain the following existence result, whose proof can be found in Appendix A.
Theorem 2.2 (Existence of minimizers for the signed distance loss).
@@ 3 Implementation
We require some additional ingredients for both steps. For the heat method (cf. section 3.1), we describe the quadrature rules for numerically approximating the integrals incorporating the point cloud density. Additionally, we introduce a "far field" heat solution that allows us to use a very small time step while still obtaining non-zero heat throughout the entire domain. For the SDF computation (cf. section 3.1.2), we explain the definition of the regions $\mathcal{B}^{\pm}$ that generalize the heat method to unoriented point clouds and their incorporation in the loss function using a smoothed characteristic function.
Figure 5: Left: non-uniform input point cloud on a square (qualitative visualization), middle: heat solution $u^{\tau}$ in two dimensions using uniform weights for the mean (hyper)surface integral 12 and right: heat solution using the locally adapted weights 11.
@@ 3.1.1 Quadrature & Point Clouds
To evaluate the loss functions, we need of course, to numerically approximate the involved integrals. To this end, we scale all the examples to $[-1,1]^{3}$ (with a distance of at least $0.2$ to the boundary of $\Omega$). Then, in every iteration of the descent algorithm, we uniformly sample points from $\Omega$ and approximate the integral by the average of the integrand’s values at these sampled points. This number of quadrature points was set to $10000$ for all experiments. The first step of the algorithm involves the evaluation of the functional $u\mapsto\fint_{\mathcal{S}}u\,\mathrm{d}a$. The surface $\mathcal{S}$ is described by a point cloud $(x_{i})_{i=1,\ldots,N}\subset\mathbb{R}^{3}$. In case the point cloud is uniformly distributed on $\mathcal{S}$, one might approximate the mean surface integral by
@@
In the general case of a non-uniform distribution of the points on the surface $\mathcal{S}$, we use the following local averaging strategy. Let $\nu\colon\mathbb{R}^{3}\to\mathbb{R}$ be a compactly supported, non-negative, smooth function, e.g., $\nu(s)=\exp((|s|^{2}-1)^{-1})$ for $|s|<1$ and 00 else. Then, for $\nu^{\epsilon}(s)=\epsilon^{-3}\nu(\epsilon^{-1}s)$ we first define the non-normalized local adaptive weights
@@
as a density rescaled approximation of the integral mean of $u$ on $\mathcal{S}$. In fig. 5, we show the heat solution $u^{\tau}$ for a non uniform input point cloud comparing uniform weights 11 with the local adaptive weights 12. It demonstrates that a proper weighting is necessary for non-uniform point clouds to obtain accurate normal directions from the heat solution.

### v2

@@ 1 Introduction
We propose to lift the heat method to the neural setting (see LABEL:fig:teaser). To that end, we define two well-posed variational problems, which on the one hand, can be optimized well with stochastic gradient descent using a standard network architecture, and on the other hand, have existence guarantees. As in the heat method, in the first step we solve for a time step of the heat flow and compute using this the gradient of the unsigned distance function, exploiting the relation between the heat kernel and the distance. Unlike in the heat method, we do not have a discretization of the domain, as we work with neural representations. Thus, to compute the gradients, we use a variational formulation for a backward Euler timestep of the heat flow [33] in the spirit of the general class of minimizing movement schemes[28], see Figure˜2 (center).
@@
In the second step, we fit the computed gradients to obtain the signed distance, taking into consideration that the input point cloud is unoriented, leading to accurate level sets in a narrow band near the surface; see Figure˜2 (right).
@@ 1.1 Related Work
The seminal paper SIREN [32] uses an eikonal loss, which combined with periodic activation functions, leads to mostly good results. As the results are very initialization dependent, DiGS [2] has proposed a geometric initialization that improves the result for unoriented point clouds, as well as an additional regularization with a divergence-based loss. StEik [37] has discussed the instabilities inherent in the eikonal equation and proposed additional regularizing losses, as well as using a quadratic neural network instead of the usual linear one. Concurrently with our work, HotSpot [35], building on the quadratic network of StEik, has proposed adding a regularizing loss that encourages the exponent of the SDF to fulfill the heat equation. This approach, like ours, incorporates the relation between the distance function and the heat equation. However, their formulation is still based on the eikonal equation, making it non-convex and more difficult to optimize, whereas we propose to solve two convex functionals successively. Furthermore, the density of the point cloud is not taken into account in their method, leading to problematic regions for non-uniform point cloud sampling.
@@
Minimizing movements schemes are a robust and effective time discretization for gradient flow type evolution problems in metric spaces [10]. In [28], deep learning-based minimizing movement schemes for a variety of PDEs of gradient flow type are proposed. In [16], a structure-preserving Eulerian neural network discretization for the numerical solving of $L^{2}$-gradient flows is presented. These schemes are based on the variational principle of the energy-dissipation law, which ensures the monotonic decay of the free energy and avoids unphysical states. We take inspiration from these machine learning approaches to properly capture the gradient flow dynamics for the heat equation in Euclidean and Riemannian spaces.
@@
Figure 3: Comparison to grid-based method [15].
Using the heat equation for computing approximate geodesic distances has been introduced to geometry processing by Crane et al. [8, 9]. This method is based on two main steps. First, diffusing heat for a short time step starting from heat concentrated at the source points. Then, the normalized gradients of the heat function are used to compute the approximate geodesic distances by solving a Poisson equation. Many generalizations and improvements have been proposed, the latest being generalization to signed distance functions (GSD) [15] by diffusing signed normals. To the best of our knowledge, heat-based distance approximations have not been used in the neural setting so far. Following the original heat method, we also diffuse a scalar function from the input sources. However, we incorporate additional constraints in the second step, which allows us to correctly orient the heat gradients. Hence, unlike GSD, we are not diffusing normals but start with unoriented point clouds to compute a signed distance field. A comparison of our neural approach with $\approx 64^{3}$ parameters with the results of the GSD method [15] for two different grid sizes is provided in the inset Figure˜3. Note that for a grid size of $64$ (middle), where the number of variables corresponds to our number of parameters, the GSD reconstruction has considerable artifacts (e.g. fingers merging). To obtain a reconstruction comparable to ours, a larger grid size is required (right).
@@
  • compute neural SDFs which are accurate not only on the surface, but also in a narrow band near the surface,
  • handle spatially varying point cloud densities,
@@ 2 Method
Throughout this paper, we consider a surface $\mathcal{S}$ which is the boundary of a Lipschitz set, contained in $1,1]^{3}$ such that the complement of this set is connected. Later, we will assume additional smoothness of $\mathcal{S}$. Following the original heat method [8], our method comprises two steps. First (Section˜2.1), we compute the solution $u\colon\mathbb{R}^{3}\to\mathbb{R}$ of a small time step of the heat equation with the mean surface measure–reflecting the density of the point cloud–as initial data. This measure is approximated via a weighted sum of point measures of a not necessarily uniformly distributed input point cloud. In the second step (Section˜2.2), we compute the SDF $\phi\colon\mathbb{R}^{3}\to\mathbb{R}$, by requiring that (1) $|\phi|$ is small on the set of input points, (2) the direction of $\nabla\phi$ is aligned with the direction of $\nabla u$. To allow unoriented point clouds as input, we automatically find regions $\mathcal{B}^{+}$ and $\mathcal{B}^{-}$ on the two sides of the cloud, and encourage $\phi$ to have the correct sign in these regions. Both $u$ and $\phi$ are parameterized with a neural network (see Section˜3 for the details). The regions $\mathcal{B}^{\pm}$ are computed using a simple deterministic algorithm.
2.1 Unoriented Point Cloud to Unoriented Normals
In the first step, we use that the normalized gradients of a short-time solution of the heat equation on $\Omega$, with the mean surface area density of a smooth surface as initial data, approximate the gradient of the distance function ${\mathrm{dist}\,}(\cdot,\mathcal{S})$ in the vicinity of the surface (cf. Crane et al [8] for the case on manifolds). We restrict here to a computational domain $\Omega\coloneqq(-1.2,1.2)^{3}$ where the surface $\mathcal{S}$ is scaled to fit into $1,1]^{3}$. In the simple case, for initial data $u^{0}\in L^{2}(\Omega)$ a single timestep $\tau$ of the heat equation
@@ 2.1 Unoriented Point Cloud to Unoriented Normals
with $u^{\tau}\approx u(\tau,\cdot)$. We want to parametrize the solution $u$ using a neural network. Here, we prefer a variational approach to compute the solution to ˜2 over a least-squares loss for the residual of ˜1, as this would decrease the numerical stability. To this end, we introduce the energy
${\mathcal{E}_{\mathrm{MM}}}(u)\coloneqq\int_{\Omega}(u-u^{0})^{2}+\tau|\nabla u|^{2}\,\mathrm{d}x\,.$ (3)
Indeed, the associated Euler-Lagrange equation characterizing the minimizer $u^{\tau}\coloneqq\operatorname*{arg\,min}_{u}{\mathcal{E}_{\mathrm{MM}}}(u)$ is given by
$0=\partial_{u}{\mathcal{E}_{\mathrm{MM}}}[u^{\tau}](\vartheta)=2\int_{\Omega}(u^{\tau}-u^{0})\vartheta+\tau\nabla u^{\tau}\cdot\nabla\vartheta\,\mathrm{d}x$
for all test functions $\vartheta\in H^{1}(\Omega)$, which is the weak formulation of (2). Here, $H^{1}(\Omega)$ is the space of $L^{2}$ functions on $\Omega$ with weak derivatives also in $L^{2}$. We remark that this corresponds to a simple example of a minimizing movements approach for the solution of ˜1 [10] (cf. [28] in the context of neural networks).
The case of interest in this paper is the mean two-dimensional surface measure as initial data $u^{0}$, which is a measure and not an element of $L^{2}(\Omega)$. The application of this measure to a function $\phi$ is defined as $\fint_{\mathcal{S}}\phi\,\mathrm{d}a\coloneqq\left(\int_{\mathcal{S}}\,\mathrm{d}a\right)^{-1}\int_{\mathcal{S}}\phi\,\mathrm{d}a$. When $\mathcal{S}$ is the boundary of a Lipschitz set, this measure is well-defined on $L^{2}$ functions on $\mathcal{S}$ and thus by the trace theorem on $H^{1}(\Omega)$.
In order to apply a similar variational approach for the solution of ˜2 in case of a measure as initial data we need to modify the energy $\mathcal{E}_{\mathrm{MM}}$}. We expand $\int_{\Omega}(u-u^{0})^{2}\,\mathrm{d}x$ and skip the term $\int_{\Omega}(u^{0})^{2}\,\mathrm{d}x$, which is irrelevant for the minimization since it does not depend on the variable $u$. Thus, we consider the modified minimizing movement energy
${\mathcal{E}^{\mathrm{mod}}_{\mathrm{MM}}}(u)\coloneqq\int_{\Omega}u^{2}\,\mathrm{d}x-2\fint_{\mathcal{S}}u\,\mathrm{d}a+\tau\int_{\Omega}|\nabla u|^{2}\,\mathrm{d}x.$ (4)
Defining as above $u^{\tau}\coloneqq\operatorname*{arg\,min}_{u}{\mathcal{E}^{\mathrm{mod}}_{\mathrm{MM}}}(u)$, the associated Euler-Lagrange equation for $u^{\tau}$ is given by
@@
for all $\vartheta\in H^{1}(\Omega)$, which is the adapted weak form of (2). The following proposition provides the existence of the discrete heat solution $u^{\tau}$.
@@
Under the above assumptions there exists a unique minimizer $u^{\tau}$ on $H^{1}(\Omega)$, which solves $0=\partial_{u}{\mathcal{E}^{\mathrm{mod}}_{\mathrm{MM}}}(u)$.
The proof can be found in Appendix˜A.
@@
which we assume to be well-defined almost everywhere, yields for smooth surfaces $\mathcal{S}$ a good approximation for the gradient of the distance function, i.e. the unoriented normal field of $\mathcal{S}$, which is well-defined up to the medial axis. However, note that this gradient is consistent only in the near field of the surface and corresponds to the unsigned distance and not the signed we actually would like to compute (Figure˜4).
@@
2.2 Unoriented Normals to SDF
In the second step, we assume that we are given an approximate, unoriented normal field $n^{\tau}$ and aim for the SDF $\phi$ of the surface. To this end, we pick up the ansatz by Crane et al. [8] and consider a modification of the second step of their heat method. We construct an SDF approximation from $n^{\tau}$ with $\phi<0$ inside the surface, where the gradient should point in the opposite direction of $n^{\tau}$, and conversely, $\phi>0$ outside the surface with the gradient pointing in the same direction as $n^{\tau}$ (Figure˜5). To this end, we seek a minimizer of the normal fitting objective
@@ 2.2 Unoriented Normals to SDF
where $\phi<0]=\{x\in\Omega\mid\phi(x)<0\}$ and $\phi\geq 0]=\{x\in\Omega\mid\phi(x)\geq 0\}$. Here, to have a decomposition of the domain, we arbitrarily included the set $\{\phi=0\}$ in the second term, which does not make a difference since this set is expected to be lower dimensional. Furthermore, we want the zero-level set of $\phi$ to coincide with the given surface, i.e. we take into account the surface fitting objective
@@
The sum ${\mathcal{E}_{n}}+{\mathcal{E}_{\mathrm{fit}}}$ has at least four minima: the signed distance, the unsigned distance, and their respective negatives. Hence, we need an additional term to favor the SDF as the desired solution. For this, we assume to be given sets $\mathcal{B}^{-}$ on the inside of $\mathcal{S}$ and $\mathcal{B}^{+}$ on the outside, cf. Section˜3.2.1 on how to generate these sets automatically. Now, we define the orientation objective
@@
which promotes that $\phi\leq 0$ in $\mathcal{B}^{-}$ and $\phi\geq 0$ in $\mathcal{B}^{+}$. Here, we denote by $\chi_{[\phi>0]}$ and $\chi_{[\phi<0]}$ the characteristic functions of the respective sets. That is, $\chi_{[\phi>0]}(x)=1$ if $\phi(x)>0$ and 0 otherwise; similarly for $\chi_{[\phi<0]}$. Finally, our overall objective for computing the SDF becomes
@@
In fact, it will turn out that, in our descent scheme, $\mathcal{E}_{\mathcal{B}}$} vanishes after a few descent iterations, which allows us to choose $\lambda_{\mathcal{B}}=1$. We obtain the following existence result, whose proof can be found in Appendix˜A.
Theorem 2.2 (Existence of minimizers for $\mathcal{E}_{\mathrm{SDF}}$}).
@@ 3 Implementation
We require some additional ingredients for both steps. For the heat method (cf. Section˜3.1), we describe the quadrature rules for numerically approximating the integrals incorporating the point cloud density. Additionally, we introduce a "far field" heat solution that allows us to use a very small time step while still obtaining non-zero heat throughout the entire domain. For the SDF computation (cf. Section˜3.2), we explain the definition of the regions $\mathcal{B}^{\pm}$ that generalize the heat method to unoriented point clouds and their incorporation in the loss function using a smoothed characteristic function.
Figure 6: Left: non-uniform input point cloud on a square (qualitative visualization), middle: heat solution $u^{\tau}$ in two dimensions using uniform weights for the mean (hyper)surface integral ˜12 and right: heat solution using the locally adapted weights ˜11.
@@ 3.1.1 Quadrature & Point Clouds
To evaluate the loss functions, we need of course, to numerically approximate the involved integrals. To this end, we scale all the examples to $1,1]^{3}$ (with a distance of at least $0.2$ to the boundary of $\Omega$). Then, in every iteration of the descent algorithm, we uniformly sample points from $\Omega$ and approximate the integral by the average of the integrand’s values at these sampled points. This number of quadrature points was set to $10000$ for all experiments. The first step of the algorithm involves the evaluation of the functional $u\mapsto\fint_{\mathcal{S}}u\,\mathrm{d}a$. The surface $\mathcal{S}$ is described by a point cloud $(x_{i})_{i=1,\ldots,N}\subset\mathbb{R}^{3}$. In case the point cloud is uniformly distributed on $\mathcal{S}$, one might approximate the mean surface integral by
@@
In the general case of a non-uniform distribution of the points on the surface $\mathcal{S}$, we use the following local averaging strategy. Let $\nu\colon\mathbb{R}^{3}\to\mathbb{R}$ be a compactly supported, non-negative, smooth function, e.g., $\nu(s)=\exp((|s|^{2}-1)^{-1})$ for $|s|<1$ and 0 else. Then, for $\nu^{\epsilon}(s)=\epsilon^{-3}\nu(\epsilon^{-1}s)$ we first define the non-normalized local adaptive weights
@@
as a density rescaled approximation of the integral mean of $u$ on $\mathcal{S}$. In practice, $\epsilon$ is chosen such that every $\epsilon$-ball centered at any point contains at least 12 neighboring points. In Figure˜6, we show the heat solution $u^{\tau}$ for a non uniform input point cloud comparing uniform weights ˜11 with the local adaptive weights ˜12. It demonstrates that a proper weighting is necessary for non-uniform point clouds to obtain accurate normal directions from the heat solution.